# AI-Reporter: A Path to a New Genre of Scientific Communication

### v1

@@ Abstract
The AI-Reporter represents a paradigmatic shift in scientific publication practice. This document demonstrates through a concrete case study how our system transforms academic presentations into publication-ready chapters—in less than three minutes. Using Arno Simons’ lecture on Large Language Models from the “Large Language Models for the History, Philosophy, and Sociology of Science” workshop (NEPI) as an example, we show how technological innovation bridges the gap between ephemeral presentation and permanent scientific documentation.
@@ 2 Case Discussion: Opening Presentation by Arno Simons
To demonstrate the functionality of our system, we selected the opening presentation by Arno Simons on “Large Language Models for the History, Philosophy, and Sociology of Science” from the NEPI workshop at TU Berlin. We are grateful to Dr. Simons for his permission to use his contribution for this detailed documentation of the processing workflow. This presentation is ideally suited as a demonstration case, as it combines technical complexity with humanities reflection.

### v2

@@ Abstract
The AI-Reporter represents a paradigmatic shift in scientific publication practice. This document demonstrates through a concrete case study how our system transforms academic presentations into publication-ready chapters—in less than three minutes. Using Arno Simons’ lecture on Large Language Models (LLMs) from the “Large Language Models for the History, Philosophy, and Sociology of Science” workshop as an example, we show how technological innovation bridges the gap between ephemeral presentation and permanent scientific documentation.
@@ 2 Case Discussion: Opening Presentation by Arno Simons
To demonstrate the functionality of our system, we selected the opening presentation by Arno Simons on “Large Language Models for the History, Philosophy, and Sociology of Science” from the NEPI workshop at TU Berlin. We are grateful to Arno Simons for his permission to use his contribution for this detailed documentation of the processing workflow. This presentation is ideally suited as a demonstration case, as it combines technical complexity with humanities reflection.
@@ 9.1 Future Developments
10 Acknowledgments
Acknowledgment: I thank Adrian Wüthrich, Arno Simons, and Michael Zichert for valuable discussions and feedback during the development process of the AI reporter in the context of the mentioned workshop. Financial support by the ERC Grant project “Network Epistemology in Practice (NEPI, PI: Adrian Wüthrich)” is gratefully acknowledged. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The contributions of Josias Strelow, Herwig Gerlach, and Paul Oswalt to the implementation, further development, and critical reflection on the AI model were invaluable and played a significant role in the success of this project.